# José Luis Blanco-Claraco

### v1

@@ 4.9 Newer College dataset
Our LO method has been evaluated against the SOTA method KISS-ICP (Vizzo et al. (2023)), with quantitative results summarized in Table 9, where it can be seen that our system has a better accuracy in 11 out of the 14 sequences, 8 of them with ATE reductions larger than 50%. More over, KISS-ICP diverges in 3 of the hardest sequences (“Maths Hard” (MH), “Underground Hard” (UH), “Stairs” (ST)) while MOLA-LO achieves ATE better than 0.12 m in all of them. It is worth mentioning that none of the two benchmarked LO methods make use of the IMU data. Other works, like Pfreundschuh et al. (2024), have recently reported even better accuracy in novel methods exploiting the intensity channel of LiDAR points together with inertial measurements, but for the sake of fair benchmarking we limit the present comparison to methods relying solely on 3D points.
Trajectories obtained by our LO method, and by the loop-closure post-processing stage, are illustrated in Figures 1920, together with ground truth. In order to understand why the sequences labeled as “hard” are harder to cope with, we show the translational ($x$,$y$,$z$) and orientation (yaw, pitch, roll) components of the trajectory estimated for one such sequence (“Maths hard”), along with ground truth, in Figure 21. Observe how, in time steps corresponding to $t\approx 140~{}s$, $t\approx 180~{}s$, and $t\approx 225~{}s$, the sensor is rotated at a high speed in different combinations of orientations and translations. The introduction of the joint estimation of the velocity vector within the ICP loop itself (see §3.7.3) has revealed essential to cope with these sudden disruptions of the constant velocity assumption.
@@ 4.10 UAL VLP-16 campus dataset
Finally, the proposed LO system, the full SLAM system including loop closure, and the SOTA method KISS-ICP (Vizzo et al. (2023)) have been benchmarked in the automotive dataset reported in Blanco-Claraco et al. (2019), where an electric vehicle is driven around the University of Almeria campus while grabbing scans from a Velodyne VLP-16. Ground truth trajectory is available from RTK GNSS 3D positioning. Estimated trajectories are shown in Figure 22, where it can be seen that, as already explained above, the largest part of the obtained ATE values is caused by accumulation of errors in the vehicle altitude above the initial ground level. Such error is drastically reduced by handling loop closures, as can be seen in Figure 22(b). Regarding a quantitative analysis of errors, they are available in Table 10, with our system exhibiting $\sim 30\%$ less ATE than the baseline method (KISS-ICP), while still being able to run LO $\sim 6$ times faster than the sensor rate.
@@
Figure 22: Estimated trajectories for the proposed LiDAR odometry system (”MOLA-LO”) applied to the UAL VLP-16 Campus dataset, compared to ground truth. Trajectories denoted as “MOLA-LO + LC” include the post-processing loop-closure. See discussion in §4.10.
@@ 5 Qualitative demonstrations
After experimentally measuring the accuracy of our proposal for several public 3D LiDAR datasets with known ground truth, we now provide evidence of other qualitative features illustrating different claims of this work.
@@
The flexibility of our framework to cope with different sensors is here illustrated with an alternative configuration, dubbed lidar-2d, suitable for mapping with 2D range finders. While all experiments shown so far comprise building a 3D point cloud local map from 3D scans, just two configuration changes enable the architecture in Figure 3 to build maps from 2D sensors. First, the local map generator (block #5 in Figure 3) is set to create an occupancy voxel map. A plain 2D grid map would also work but, as discussed in §3.3, the sparse data structure of voxel maps avoids the need for frequent memory reallocations as the robot explores. Secondly, the observation pipeline (block #3 in Figure 3) is simplified to that shown in Figure 23, that is, there is no need to down-sample the raw sensor points twice (for registration and for map update) due to the reduced size of 2D scans.
Figure 23: Contents of the “observation pipeline” blocks in Figure 3 for the “2D LiDAR” LO system configuration. Refer to discussion on pipelines in general in §3.4 and to §5.1 for this particular pipeline.
This alternative configuration has been validated with the datasets exposed below, all of them from wheeled robots equipped with encoders from which incremental odometry is taken as an input for the kinematic state prediction module (§3.9). First, we applied the LO system to the Freiburg building 079 (fr079) dataset (Howard and Roy (2003)), recorded by Cyrill Stachniss in 2010 with a Pioneer2 mobile robot equipped with a SICK LMS range finder. The resulting voxel map (which is only populated at one fixed height) is illustrated in Figure 24(a). Our method took an average of 10.2 ms to process each scan. Note that loop closures were not required here.
@@
Figure 24: Results for the 2D LiDAR configuration. Refer to discussion in §5.1.
@@
Figure 25: Illustrative results for metric map georeferenciation for sequences estimated by the SLAM solution and automatically georeferenced as described in §5.2 from consumer-grade GNSS data (Screenshot from Google Earth v7.3.6.9326, images by: Maxar Technologies, 2024; Airbus, 2024).)
@@
Figure 26: Final 3D maps from datasets collected in Almería forests with a LiDAR backpack. See discussion in §5.3.
@@ 5.2 Georeferencing metric maps
Finding the geodetic coordinates of metric maps is possible from a low-cost GNSS receiver given observations enough, following the method described in §3.12. In this section we show experimental results in georeferencing maps, based on the MulRan dataset (see §4.1) since it includes both, accurate ground truth, and readings from a low-cost GNSS device.
@@
Figure 27: Configuration for building metric maps of arbitrary types from view-based maps using the application sm2mm. The type of the final map is solely determined by the “Generator” block on the bottom, without any other required change in the rest. Refer to discussion in §5.4.
@@
Figure 28: Example final maps built from the Voxgraph dataset. The voxel map in (a) has been visualized as one point per voxel to favor a clearer visualization. Refer to discussion in §5.4.
@@ 5.4 Building different metric maps
One of the claims of this work is that different robotic tasks are better performed using different map representations, thus the mapping framework should be able to provide such flexibility. To illustrate this feature, we show next how easy is to create metric maps of different types, apart of the commonly-used point clouds. Focusing on the post-processing stage, once mapping is done and there is a view-based map from the LO or SLAM system, it was explained in §3.11 how such map can be converted into metric maps. In particular, Figure 8 illustrated how to build several map layers by accumulating 3D LiDAR scans into a global map.
Here, we use a modification of such pipeline, shown in Figure 27, where only one final map layer exists of a custom type which will be changed between experiments. The final map obtained from the Voxgraph dataset (§4.7) was fed into that pipeline with two map types: an occupancy voxel map, and a 2D digital elevation model (DEM). The resulting maps are illustrated in Figure 28(a)–(b). Changing a few lines in a configuration YAML file is all that it takes to build one map type or the other.
6 Localization experiments
Localization without mapping can be performed in the present framework in two ways: (i) with particle filtering (as in Blanco-Claraco et al. (2019)) using the mrpt_pf_localization software package, or (ii) directly with the LO pipeline introduced in §3.5, by disabling map updates. In both cases, the reference map can be given from converting a view-based map from a mapping session into a metric map, as described in §3.11. Robots moving slowly and equipped with wheels encoders may be good fits for the particle filtering approach. For faster vehicles and drones, the LO pipeline is probably a better option due to the more accurate initial predictions for each time step given by the kinematic state prediction module (§3.5).
@@ 6 Localization experiments
Figure 29: Results of the localization experiments performed with MulRan dataset sequences KAIST01 and KAIST03 by using as reference the map built from KAIST02. Refer to discussion in §6.
@@
In this section, we provide experimental results for localization using the MulRan dataset, which perfectly fits this task since three driving sequences exist for each location. Therefore, a metric map from the SLAM output for one such sequence (KAIST02) is built and used as reference map for localizing the other two sequences of the same location (KAIST01 and KAIST03). The reference global map was built with the metric map building pipeline in Figure 7, using a point cloud as global map. In particular, from all point cloud maps described in §3.3, we selected the one with hashed voxels as underlying data structure, with a 2 m resolution and a maximum of 20 points per voxel. Localization errors have been estimated using evo_ape (Grupp (2017)) between the estimated trajectories and ground truth, as illustrated in Figure 29. In this case, the initial approximated pose was given manually for each sequence, with an error of several meters, hence the error peak at the beginning of each sequence. Automatically re-localizing from GNSS readings is possible as a more practical alternative, but at the time of writing, this feature is only implemented in the particle filter-based solution. The localization RMSE with respect to ground truth in these two experiments were 2.21 m and 2.11 m, respectively. Note, however, that virtually all this error is attributable to accumulated residual drift in the map used as reference, since 2.13 m is precisely the RMSE ATE of the full SLAM solution for sequence KAIST02 (refer to row “MOLA-LO+LC” in Table 2). Basically, localization accuracy is limited by how accurate the reference map is. Note that the tested sequences include dynamic objects (e.g. vehicle overtaking), starts and stops due to traffic, etc.
@@ 7 Conclusions
This paper introduced a whole framework that aims at filling a gap in the robotics community’s needs for flexible map building and editing from 3D LiDAR. It has been shown how the proposed LO and SLAM systems compare well or outperform other SOTA methods regarding estimated trajectory accuracy and robustness against divergence. Most parts of the framework are available as open-source software. The present work leaves plenty of research topics open for future works: (i) benchmarking different metric map data structures to find out which ones suit best to each problem (e.g. real-time LO vs. localization without mapping), (ii) designing new pipeline blocks for smarter sampling of point clouds to achieve both, faster and more accurate LO, and (iii) alternative pipeline designs suitable for efficient depth camera odometry, to name a few of them.
Appendix A Software ecosystem
This appendix enumerates all software packages and libraries involved in the systems and experiments described in this paper. Part of these contributions are a follow-up of the former work Blanco (2019), which can be checkout out for details about the MOLA architecture, life cycle of modules, inter-thread communication, etc.
The whole system has been purposely designed to be fully compatible with ROS 2 (Macenski et al. (2022)), its reference build tools (i.e. colcon, ament), and its data set format (rosbags2). Compatibility with ROS 1 data sets is possible via conversion to either rosbags2 or MRPT’s own format (RawLog). Next follows a list of software packages and their most relevant contents in the context of this work:
  • MRPT (Mobile Robot Programming Toolkit): A set of C++ libraries and applications covering data structures and common algorithms usually found in mobile robotics and computer vision. All raw sensor observations in this work are stored as classes from the mrpt-obs library. Cross-platform and OS-independent binary serialization, key for persistence of maps, datasets, etc. is provided by the mrpt-serialization library. View-based maps are implemented as simple-maps, defined in the mrpt-maps library together with many of the metric map types defined in §3.3.
  • mp2p_icp (Multi Primitive-to-Primitive ICP): C++ libraries and applications related to the multi-layer metric map structure (§3.3), metric map update pipelines (§3.4), and ICP. It also provides important tools like: sm2mm (“simple map to metric map”) implementing the process described in §3.11; a graphical interface to inspect ICP log, enabling analyzing the internals of each ICP iteration during mapping; the program mm-viewer to inspect multi-layer metric maps, etc.
  • MOLA: Core packages for data-set sources (§3.2), additional definitions of metric map types (mola_metric_maps), a bridge between MOLA modules and ROS 2 (mola_bridge_ros2), etc.
  • mola_lidar_odometry: The main piece implementing the LO architecture described in §3.5.
  • mola_sm_loop_closure: This package implements georeferenciation (§3.12) and loop closure (§3.13).
  • mrpt_navigation: A set of ROS 2 packages, including: mrpt_map_server for publishing multi-layer metric maps (§3.3), and mrpt_pf_localization performing particle filter-based localization from such maps.

### v2

@@ 4.9 Newer College dataset
Our LO method has been evaluated against the SOTA method KISS-ICP (Vizzo et al., 2023), with quantitative results summarized in Table 9, where it can be seen that our system has a better accuracy in 11 out of the 14 sequences, 8 of them with ATE reductions larger than 50%. More over, KISS-ICP diverges in 3 of the hardest sequences (“Maths Hard” (MH), “Underground Hard” (UH), “Stairs” (ST)) while MOLA-LO achieves ATE better than 0.12 m in all of them. It is worth mentioning that none of the two benchmarked LO methods make use of the IMU data. Other works, like Pfreundschuh et al. (2024), have recently reported even better accuracy in novel methods exploiting the intensity channel of LiDAR points together with inertial measurements, but for the sake of fair benchmarking we limit the present comparison to methods relying solely on 3D points.
Trajectories obtained by our LO method, and by the loop-closure post-processing stage, are illustrated in Figures 2223, together with ground truth. In order to understand why the sequences labeled as “hard” are harder to cope with, we show the translational ($x$,$y$,$z$) and orientation (yaw, pitch, roll) components of the trajectory estimated for one such sequence (“Maths hard”), along with ground truth, in Figure 24. Observe how, in time steps corresponding to $t\approx 140~{}s$, $t\approx 180~{}s$, and $t\approx 225~{}s$, the sensor is rotated at a high speed in different combinations of orientations and translations. The introduction of the joint estimation of the velocity vector within the ICP loop itself (see Section 3.6.3) has revealed essential to cope with these sudden disruptions of the constant velocity assumption.
@@ 4.10 UAL VLP-16 campus dataset
Finally, the proposed LO system, the full SLAM system including loop closure, and the SOTA method KISS-ICP (Vizzo et al., 2023) have been benchmarked in the automotive dataset reported in Blanco-Claraco et al. (2019), where an electric vehicle is driven around the University of Almeria campus while grabbing scans from a Velodyne VLP-16. Ground truth trajectory is available from RTK GNSS 3D positioning. Estimated trajectories are shown in Figure 25, where it can be seen that, as already explained above, the largest part of the obtained ATE values is caused by accumulation of errors in the vehicle altitude above the initial ground level. Such error is drastically reduced by handling loop closures, as can be seen in Figure 25(b). Regarding a quantitative analysis of errors, they are available in Table 10, with our system exhibiting $\sim 30\%$ less ATE than the baseline method (KISS-ICP), while still being able to run LO $\sim 6$ times faster than the sensor rate.
@@
Figure 25: Estimated trajectories for the proposed LiDAR odometry system (”MOLA-LO”) applied to the UAL VLP-16 Campus dataset, compared to ground truth. Trajectories denoted as “MOLA-LO + LC” include the post-processing loop-closure. See discussion in Section 4.10.
@@ 5 Qualitative demonstrations
After experimentally measuring the accuracy of our proposal for several public 3D LiDAR datasets with known ground truth, we now provide evidence of other qualitative features illustrating different claims of this work. Further claims will be also analyzed later on in a quantitative way in Section 7 with ablation studies.
@@
The flexibility of our framework to cope with different sensors is here illustrated with an alternative configuration, dubbed lidar-2d, suitable for mapping with 2D range finders. While all experiments shown so far comprise building a 3D point cloud local map from 3D scans, just two configuration changes enable the architecture in Figure 3 to build maps from 2D sensors. First, the local map generator (block #5 in Figure 3) is set to create an occupancy voxel map. A plain 2D grid map would also work but, as discussed in Section 3.2, the sparse data structure of voxel maps avoids the need for frequent memory reallocations as the robot explores. Secondly, the observation pipeline (block #3 in Figure 3) is simplified to that shown in Figure 26, that is, there is no need to down-sample the raw sensor points twice (for registration and for map update) due to the reduced size of 2D scans.
Figure 26: Contents of the “observation pipeline” blocks in Figure 3 for the “2D LiDAR” LO system configuration. Refer to discussion on pipelines in general in Section 3.3 and to Section 5.1 for this particular pipeline.
This alternative configuration has been validated with the datasets exposed below, all of them from wheeled robots equipped with encoders from which incremental odometry is taken as an input for the kinematic state prediction module (Section 3.8). First, we applied the LO system to the Freiburg building 079 (fr079) dataset (Howard and Roy, 2003), recorded by Cyrill Stachniss in 2010 with a Pioneer2 mobile robot equipped with a SICK LMS range finder. The resulting voxel map (which is only populated at one fixed height) is illustrated in Figure 27(a). Our method took an average of 10.2 ms to process each scan. Note that loop closures were not required here.
@@
Figure 27: Results for the 2D LiDAR configuration. Refer to discussion in Section 5.1.
@@
Figure 28: Illustrative results for metric map georeferenciation for sequences estimated by the SLAM solution and automatically georeferenced as described in Section 5.2 from consumer-grade GNSS data (Screenshot from Google Earth v7.3.6.9326, images by: Maxar Technologies, 2024; Airbus, 2024).)
@@
Figure 29: Final 3D maps from datasets collected in Almería forests with a LiDAR backpack. See discussion in Section 5.3.
@@ 5.2 Georeferencing metric maps
Finding the geodetic coordinates of metric maps is possible from a low-cost GNSS receiver given observations enough, following the method described in Section 3.11. In this section we show experimental results in georeferencing maps, based on the MulRan dataset (see Section 4.1) since it includes both, accurate ground truth, and readings from a low-cost GNSS device.
@@
Figure 30: Configuration for building metric maps of arbitrary types from view-based maps using the application sm2mm. The type of the final map is solely determined by the “Generator” block on the bottom, without any other required change in the rest. Refer to discussion in Section 5.4.
@@
Figure 31: Example final maps built from the Voxgraph dataset. The voxel map in (a) has been visualized as one point per voxel to favor a clearer visualization. Refer to discussion in Section 5.4.
@@ 5.4 Building different metric maps
One of the claims of this work is that different robotic tasks are better performed using different map representations, thus the mapping framework should be able to provide such flexibility. To illustrate this feature, we show next how easy is to create metric maps of different types, apart of the commonly-used point clouds. Focusing on the post-processing stage, once mapping is done and there is a view-based map from the LO or SLAM system, it was explained in Section 3.10 how such map can be converted into metric maps. In particular, Figure 8 illustrated how to build several map layers by accumulating 3D LiDAR scans into a global map.
Here, we use a modification of such pipeline, shown in Figure 30, where only one final map layer exists of a custom type which will be changed between experiments. The final map obtained from the Voxgraph dataset (Section 4.7) was fed into that pipeline with two map types: an occupancy voxel map, and a 2D digital elevation model (DEM). The resulting maps are illustrated in Figure 31(a)–(b). Changing a few lines in a configuration YAML file is all that it takes to build one map type or the other.
@@
Figure 32: Results of the localization experiments performed with MulRan dataset sequences KAIST01 and KAIST03 by using as reference the map built from KAIST02. Refer to discussion in Section 6.
6 Localization experiments
Localization without mapping can be performed in the present framework in two ways: (i) with particle filtering (as in Blanco-Claraco et al. (2019)) using the mrpt_pf_localization software package, or (ii) directly with the LO pipeline introduced in Section 3.4, by disabling map updates. In both cases, the reference map can be given from converting a view-based map from a mapping session into a metric map, as described in Section 3.10. Robots moving slowly and equipped with wheels encoders may be good fits for the particle filtering approach. For faster vehicles and drones, the LO pipeline is probably a better option due to the more accurate initial predictions for each time step given by the kinematic state prediction module (Section 3.4).
@@ 6 Localization experiments
In this section, we provide experimental results for localization using the MulRan dataset, which perfectly fits this task since three driving sequences exist for each location. Therefore, a metric map from the SLAM output for one such sequence (KAIST02) is built and used as reference map for localizing the other two sequences of the same location (KAIST01 and KAIST03). The reference global map was built with the metric map building pipeline in Figure 7, using a point cloud as global map. In particular, from all point cloud maps described in Section 3.2, we selected the one with hashed voxels as underlying data structure, with a 2 m resolution and a maximum of 20 points per voxel. Localization errors have been estimated using evo_ape (Grupp, 2017) between the estimated trajectories and ground truth, as illustrated in Figure 32. In this case, the initial approximated pose was given manually for each sequence, with an error of several meters, hence the error peak at the beginning of each sequence. Automatically re-localizing from GNSS readings is possible as a more practical alternative, but at the time of writing, this feature is only implemented in the particle filter-based solution. The localization RMSE with respect to ground truth in these two experiments were 2.21 m and 2.11 m, respectively. Note, however, that virtually all this error is attributable to accumulated residual drift in the map used as reference, since 2.13 m is precisely the RMSE ATE of the full SLAM solution for sequence KAIST02 (refer to row “MOLA-LO+LC” in Table 2). Basically, localization accuracy is limited by how accurate the reference map is. Note that the tested sequences include dynamic objects (e.g. vehicle overtaking), starts and stops due to traffic, etc.
7 Ablation studies
(a) Mulran (automotive)
(b) Newer College Dataset (handheld)
Figure 33: Results of the ablation experiments to evaluate the proposed selective map update feature, showing how it becomes relevant for datasets with irregular motion profiles (e.g. handheld sensors). Refer to discussion in Section 7.2.
7.1 Scan deskewing
First, we want to validate that scan deskewing by means of trajectory interpolation on SE(3) per Eq. (2) is good enough and effectively leads to more accurate results. As a benchmark, we tested our LiDAR odometry system (without loop closure) on the UAL campus dataset in Blanco-Claraco et al. (2019) with and without deskewing the input scans, obtaining the results in Table 10. The obtained ATE values with respect to ground truth with and without deskewing are 6.65 m and 7.97 m, respectively. Therefore, undistorting the scans reduced the ATE in a 16.6%, despite the fact that driving speed was relatively slow in this dataset, an average of $\sim 3~{}m/s$ (10.8 km/h or 6.7 mph). For comparison, enabling deskewing in KISS-ICP also reduced the ATE from 13.21 m to 10.06 m, a 23.8% improvement. Therefore, we can conclude that deskewing is absolutely a must for accurate localization, even at reduced speeds.
7.2 Local map updates
As mentioned in Section 3.4, our system differs from others in that we do not always update the local map for every single time step. The reason behind this decision is that small localization errors and imperfect scan de-skewing, specially during abrupt angular accelerations would lead to map insertion of misplaced points which would then serve as (incorrect) matches for successive scans, typically leading to unbounded error growth in localization. We postulate that such misplaced points are more likely in high-dynamics datasets than in those with smoother trajectories; e.g. handheld vs. automotive. This effect is expected when input scans are subsampled before registration against a local metric maps that do not fuse information from different scans, such as the point clouds used in our default system configuration and in many recent LiDAR odometry approaches. Probabilistic map representations or the 3D-NDT map described in Section 3.2 are expected to be more stable against frequent updates.
To support our hypothesis, we performed the following ablation experiment: we compared the trajectory error with and without the proposed selective map update in an automotive (Mulran) and a handheld (Newer College) dataset. The quantitative results were already reported in Table 2 and Table 9, respectively, but for the sake of a clearer visual comparison we have represented them in Figure 33. As can be observed, updating the map for every single time step is not problematic for the automotive dataset, with trajectory errors slightly better in some sequences and worse on others, but not leading to drastic differences. In turn, not using our proposed selective map update strategy for the Newer College dataset leads to a dramatic increase of (at least) $50\%$ for the ATE in 11 out of 14 sequences, hence supporting the introduction of our selective update strategy for robustness against high-dynamic motion profiles.
Figure 34: Evaluation of the impact of the tightly-coupled estimation of velocity vectors as part of the ablation study in Section 7.3 for the “Underground hard” sequence of the Newer College Dataset. The picture shows a bird-eye view of the estimated trajectories. In the legend, “MOLA-LO” refers to the default system as introduced in the manuscript, while “MOLA-LO without velocity estimation” is a version where tightly-coupled estimation of this vector is disabled.
7.3 Tightly-coupled estimation of vehicle velocities
Next, we want to evaluate the impact of the tightly-coupled estimation of linear and angular velocities within the optimization loop (refer to Section 3.6.3). As baseline, we use the proposal of keeping the velocity vectors as estimated from the last time step, as done in the state-of-the-art method KISS-ICP Vizzo et al. (2023). Comparison will be done using two Newer College Dataset sequences with high dynamics: “Maths hard” (MH) and “Underground hard” (UH). The trajectory ATE for MH increases from 0.11 m using the default system configuration to 0.15 m if the feature at test is disabled. A more drastic increase of the error occurs for the UH sequence, where an ATE of 0.086 m becomes 2.79 m. By visualizing the estimated paths in detail, shown in Figure 34, it can be seen that the error comes from two sources: divergence of the estimation at multiple points, and jerky trajectories for the rest. These experiments demonstrate that this feature alone has a great potential to increase the stability of LiDAR-only odometry methods.
7.4 Horn’s method
As mentioned in Section 3.6.2, the closed-form Horn’s solution to pointwise pairings would seem to be an ideal alternative to iterative methods such as Gauss-Newton due to its simplicity and efficiency. However, as we demonstrate in this section, there are two problems that render iterative non-linear methods as better practical solutions. We have compared the accuracy of our default LiDAR odometry system with another version where the Gauss-Newton optimizer is replaced with Horn’s closed form solution (a change that only takes adding three lines in a YAML configuration file). The dataset on which the study has been done is KITTI, with results already shown in Table 3 in the rows “MOLA-LO (default)” and “MOLA-LO (Horn’s)”. In all sequences, excepting 01 and 04, the Horn’s method leads to worse results. Overall, the average RTE worsens from 0.55% to 0.65%. The reason for these results is the impossibility for this method to cope with outliers, something easy to integrate in non-linear optimizers. Another interesting observation is that, despite the solver itself runs $\sim 4$ times faster than Gauss-Newton, the overall time cost is $\sim 2$ times slower due to the need to spend more ICP iterations until convergence.
7.5 Loop-closure with and without GNSS
Finally, the capability of the proposed loop closure algorithm to incorporate optional GNSS readings is put at test with the Mulran dataset. Results are shown in the two last rows of Table 2. Overall, the conclusion here is that the presence of GNSS: (i) improves the global accuracy of the trajectories (reduced ATE values in the version with GNSS in 11 out of 12 sequences), and (ii) enables identifying potential loop closures that would not be found by our metric uncertainty-based hypothesis generator in the very long loops ($>20~{}km$) of Sejong sequences.
@@ 8 Conclusions
This paper introduced a whole framework that aims at filling a gap in the robotics community’s needs for flexible map building and editing from 3D LiDAR. It has been shown how the proposed LO and SLAM systems compare well or outperform other SOTA methods regarding estimated trajectory accuracy and robustness against divergence. Excepting georeferencing and loop-closure, all other components of the framework are available as open-source software. The present work leaves plenty of research topics open for future works: (i) benchmarking different metric map data structures to find out which ones suit best to each problem (e.g. real-time LO vs. localization without mapping), (ii) designing new pipeline blocks for smarter sampling of point clouds to achieve both, faster and more accurate LO, (iii) alternative pipeline designs suitable for efficient depth camera odometry, or (iv) adding support for tightly-integrated measurements from LiDAR intensity images or inertial sensors.